TheJag18:

A landscaping business owner buys a new riding mower for $9,600.00. The owner makes a $1,000 down payment and applies for an $8,600 four-year, 3.5% interest rate installment loan, with monthly payments of $192.26. What is the total cost of the mower? $8,600.00 $9,600.00 $9,228.56 $10,228.48

Phantomdex:

n = 4 x 12 = 48 (1 year = 12 months) total cost = 1,000 + 192.26 x 48 = 10,228.48 1,000 being the down payment and 192.26 being the per month.

Phantomdex:

So, you're answer is D: $10,228.48

TheJag18:

awsome :)